OpenStudy (anonymous):

Solve the system of equations x + 3y = 32 -3x + 2y = 3

OpenStudy (anonymous):

3x+9y=64 (multiplied first by 3) add two equations: 11y=67 y=67/11

OpenStudy (anonymous):

oops that's incorrect

OpenStudy (gw2011):

x + 3y = 32 Multiply this equation by 3 in order to eliminate x -3x+2y= 3 3x +9y =96 -3x + 2y = 3 11y = 99 y = 9 x + 3(9) = 32 Substituting y=9 to find x x + 27 =32 x = 5 Therefore, x=5 and y=9

OpenStudy (anonymous):

3x+9y=96** 11y=99 y=9